OpenStudy (anonymous):

For z1 = 4cis pi/2 and z2 = 3cis 3pi/2 , find z1 · z2 in rectangular form.

zepdrix (zepdrix):

Do you understand how to write these in exponential form? \[\Large\rm 4cis\frac{\pi}{2}=4e^{\mathcal i \pi/2}\]This makes the mulltiplication easier.

OpenStudy (anonymous):

i still dont understand, it gives me no answer when i put it in the calculator

zepdrix (zepdrix):

Grr sorry I'm in class >.< hard to respond right now. So if you write each z in exponential form you have, \(\Large\rm 4e^{\mathcal i \pi/2}\cdot 3e^{\mathcal i 3\pi/2}=12e^{\mathcal i 2\pi}\)

zepdrix (zepdrix):

\(\Large\rm =12(\cos2\pi+\mathcal i \sin2\pi)=12\)